OpenStudy (anonymous):

A simple harmonic motion has amplitude 6 and circular frequency of 100pi radians per second. Write a formula for the displacement d at any time t. What is the period of the simple harmonic motion?

OpenStudy (dumbsearch2):

does this help ?

OpenStudy (dumbsearch2):

http://bit.ly/11B0lsk

OpenStudy (anonymous):

it does not help at all

OpenStudy (dumbsearch2):

haha learn to laugh ;p

OpenStudy (anonymous):

SHM : x=Acos(wt)

OpenStudy (anonymous):

So here, You are give amplitude and frequency .. Do you know the relation b/w angular frequency and frequency? -- w=2 x pi x (freq) Also, Time period is the reciprocal of frequency. Now you have all the necessary information to solve it :) :D

OpenStudy (kropot72):

SHM is a sine or a cosine function each of which has a maximum and minimum value of +1 and -1 respectively. Lets choose the sine function and give it an amplitude of 6 by multiplying by 6 and an angle in radians which is the circular (angular) frequency multiplied by t (the time in seconds). Therefore we get \[f(t)=6\sin 100 \pi t\]